OpenStudy (anonymous):

Determine the Taylor series about the poin x0 for the given function. Also determine the radius of convergence of the series.

OpenStudy (anonymous):

sin x, x0=0

OpenStudy (anonymous):

determine or remember? power series for sine is well known do you have to do it by hand?

OpenStudy (anonymous):

successive derivatives are \[\cos(x)\\ -\sin(x)\\ -\cos(x)\\ \sin(x)\] and then they repeat

OpenStudy (anonymous):

put \(x=0\) and you get \[0,1,0,-1,0,1,0,-1,...\]

OpenStudy (anonymous):

making the power series \[x-\frac{x^3}{3!}+\frac{x^5}{5!}-...\]

OpenStudy (anonymous):

Thanks